# Exploring Iterative Controllable Summarization with Large Language Models

### v2

@@ 4 Controllability of LLMs in text summarization
Llama3-8B-GTE 12.63% / 3.52 11.63% / 2.53 0.00% / 1.46 11.70% / 3.26 26.40% / 3.08 10.92% / 2.26 13.18% / 3.85 14.99% / 2.80 25.56% / 0.91 -
Llama3-70B-Iter 54.82% / 8.44 37.21% / 7.47 2.70% / 3.78 49.91% / 8.05 18.40% / 6.58 54.61% / 10.42 67.44% / 12.00 49.36% / 8.24 0.00% / 0.24 -
Llama3-70B-SAI 26.55% / 6.57 18.60% / 7.81 0.00% / 1.86 24.14% / 6.52 4.80% / 5.42 2.73% / 3.81 10.85% / 4.84 5.12% / 4.39 0.00% / 0.10 -
Llama3-70B-GTE 0.21% / 3.28 0.00% / 2.83 0.00% / 1.50 0.18% / 3.22 0.00% / 1.10 0.00% / 1.61 2.32% / 3.14 0.55% / 1.90 0.00% / 0.01 -
GPT-3.5-Iter 45.18% / 9.80 60.47% / 0.00 94.59% / 0.00 49.73% / 9.80 53.60% / 0.00 80.89% / 0.00 88.37% / 0.00 76.42% / 0.00 0.00% / 0.00 -
GPT-3.5-GTE 17.56% / 3.86 51.16% / 5.00 67.57% / 4.00 23.58% / 3.90 5.60% / 4.63 44.03% / 6.62 78.29% / 7.00 43.33% / 5.95 0.00% / 0.00 -
GPT-4o-Iter 34.69% / 6.77 55.81% / 0.00 78.38% / 3.00 39.31% / 6.63 72.00% / 0.00 64.85% / 0.00 79.07% / 0.00 69.84% / 0.00 0.38% / 0.02 -
GPT-4o-SAI 35.12% / 5.50 48.84% / 15.50 62.16% / 6.00 38.03% / 6.13 60.00% / 8.79 61.09% / 9.40 78.29% / 2.00 64.90% / 8.60 0.00% / 0.04 -
GPT-4o-GTE 0.00% / 2.76 0.00% / 4.70 0.00% / 2.03 0.00% / 2.87 0.00% / 1.20 0.00% / 1.21 0.00% / 1.96 0.00% / 1.42 0.00% / 0.02 -
Table 4: The results of controllability measured on the MACSumDoc dataset. The bold denotes the best performance.
Model Extractiveness ($\downarrow$ / $\downarrow$) Length ($\downarrow$ / $\downarrow$) Topic ($\downarrow$) Speaker ($\downarrow$)
@@
5 Method: Guide-to-Explain (GTE)
Therefore, we introduce a guide-to-explain (GTE) framework to adjust various attributes via LLMs. As shown in Figure 1, our GTE framework consists of two key phase: the step-by-step attribute identification and the self-explanation guidance. We provide an step-by-step attribute identification to teach precise adjustment for incorrectly generated responses then guide LLMs to reflect by explaining the rationales behind their mistakes. Our approach allows the model to make appropriate adjustments in subsequent iterations.
@@ 5.1 Step-by-step attribute identification
We first instruct the LLM to generate an initial summary $s^{\prime}$ that reflects the specified attribute. If the LLM fails to control the attributes accurately, we provide step-by-step attribute identification (SAI) to guide the model on how to adjust the attributes. LLM may struggle to measure numerical attributes such as extractiveness or the length ratio. Therefore, we instruct the model on how to measure the attribute step-by-step so that it can revise its previously generated summary precisely.
@@ 5.2 Self-explanation guidance
After the identification phase, we provide self-explanation guidance (SEG) to the model, guiding the model to explain why it initially failed to control the attributes. This mirrors how humans solve complex problems by reviewing their mistakes to improve future responses. Building on this, in the next iteration, the document ($d$), initial instruction ($i$), and previously generated summary ($s^{\prime}$) are provided as inputs, along with SAI and SEG. Despite LLMs being known to struggle with number-related tasks Akhtar et al. (2023); Imani et al. (2023), our guidance helps the model effectively control numerical attributes by self-explaining its miscalculations before generating summaries, especially when combined with the step-by-step attribute identification phase. We introduce GTE as a framework that integrates step-by-step attribute identification and self-explanation guidance.
@@ 5.3 Overall process
By receiving [$d$; $i$, $s^{\prime}$; SAI; SEG] as input, the model first reflects on the reasons for the initial error before generating a revised summary. If the revised summary still fails to satisfy the attributes, we repeat the GTE process until the model generates an attribute-aligned summary. In Figure 1, we illustrate in detail how the GTE framework actually operates. If the initial draft fails to properly adjust an attribute, we first provide a step-by-step guide on how to measure the attribute. Then, we guide the model to explain the miscalculation and regenerate the summary. As intended, the model mimics step-by-step attribute identification, reflects on the summary based on the identification feedback, and generates an improved summary.
@@ 6 Experimental setup
We evaluate the controllability of various LLMs, including Phi-3-medium Abdin et al. (2024), Llama3 series Dubey et al. (2024), and GPT series Brown (2020); Achiam et al. (2023). To analyze model performance by size, we utilize both the 8B 111meta-llama/Meta-Llama-3-8B-Instruct and quantized 70B versions 222casperhansen/llama-3-70b-instruct-awq of Llama3, as well as GPT-3.5 and GPT-4o. We use BERTScore Zhang et al. (2020)  [^3] https://github.com/Tiiiger/bert_score to measure embedding similarity.
We used two datasets for our experiments: MACSumDoc and the MACSumDial datasets Zhang et al. (2023b), which comprise committee meeting transcripts and news contents, respectively. Both datasets are designed for mixed-attribute summarization that controls multiple attributes simultaneously. Notably, only MACSumDial has speaker attribute. Since we evaluate LLM performance on individual attributes, we use attributes separately.
@@ 7 Results and Discussions
Figure 2: The graphs show how the length changes for each iteration. The intensity of the distribution color is proportional to the number of data points, and the markers represent the average values. The red line indicates the target length, with values of 7.5, 20, and 32.5 from left to right.
We denote the naive iteration approach, which repeatedly adjusts attributes, as Iter. The strategy that only provides step-by-step attribute identification is defined as SAI, while our full guiding framework is referred to as GTE. As shown in Table 4, our GTE demonstrate remarkably lower failure rates and require fewer iterations when adjusting summaries across all attributes, including challenging numerical attributes in MACSumDoc. Surprisingly, our method reduced the failure rate to nearly 0% when applied to Llama3-70B and GPT-4o, successfully controlling both extractiveness and length within just 1–3 iterations. For smaller models such as Phi-3-medium and Llama3-8B, which initially exhibited high failure rates, our approach significantly reduced their failure rates, demonstrating its effectiveness across different model scales. In particular, for long length, the most challenging attribute, our method achieved a remarkably low failure rate of just 2.32% within an average of 3.14 iterations.
LLMs encounter more difficulties with the MACSumDial dataset (Table 5). The dataset, which is derived from QMSum Zhong et al. (2021), consists of lengthy and diverse content parliamentary and committee meetings, making it more challenging compared to the CNN-news-based MACSumDoc. Notably, length control proved to be the most challenging attribute in the MACSumDial dataset. This difficulty is likely due to the original dataset being a lengthy parliamentary corpus, making it inherently challenging to generate summaries of a specific length. While the model handled short-length summaries well, the difficulty increased significantly as the requested summary length grew. In fact, for long length, both GPT-4o-Iter and GPT-4o-SAI showed a 100% failure rate. However, our framework meaningfully improved length controllability. With GPT-4o, the average failure rate dropped below 50%. Especially, for normal-length summaries, the failure rate further reduced from over 90% to 9.82%. Regarding extractiveness, the Iter and SAI of GPT-4o exhibit relatively low iteration count since the models mostly exceed the maximum iteration. While they fail nearly 80%, our GTE demonstrates a markedly lower failure rate at 17.28% with low iterations, demonstrating the effectiveness of our framework.
@@
To analyze how attribute changes each step, we track length adjustments per iteration (Figure 2). While all methods start with a similar distribution at the initial points, GTE consistently converges within approximately three iterations, maintaining a stable length adjustment pattern across samples. In contrast, Iter and SAI show inconsistent changes across samples, leading higher variance in length changes. This demonstrates that our method enables robust attribute control with fewer iterations, regardless of the data sample. For the experiment, we utilize Llama3-70B and randomly selected 110 samples from the MACSumDoc test set.
@@
We observe that LLMs control linguistic attributes (topic and speaker) better than numerical attributes (extractiveness and length). This aligns with previous research in mathematical reasoning, where LLM struggle with numerical features Akhtar et al. (2023), highlighting a broader challenge in precisely handling numerical constraints. From the perspective of the summarization task, extractiveness and length control the structure of the summary, whereas topic and speaker control its contents. Our findings suggest that LLMs are proficient at adjusting content to align with user preferences, but they struggle to generate summaries in a desired structural format.
@@
Table 6: Quality of the controlled summaries.
@@
We evaluate the quality of summaries generated by GTE. We mainly use UniEval Zhong et al. (2022) and QuestEval Scialom et al. (2021), as they highly correlate with human judgement and assess the overall quality of the summary itself. UniEval is a multi-dimensional evaluator that assesses coherence, consistency, fluency, and relevance of summaries. QuestEval measures precision and recall by leveraging question answering framework to compare the content between the source document and the generated summary without relying on the reference summary. Table 6 shows that our method’s summaries outperform across all UniEval dimensions and QuestEval, demonstrating effective attribute control while maintaining overall summary quality. Relevance assesses how well a summary retains key information compared to the reference. While Iter and SAI generate misaligned summaries with lower relevance scores, GTE effectively aligns them, leading to a substantial gain.
Although previous studies have shown that ROUGE scores Lin (2004) are insufficient for assessing summary quality Scialom et al. (2021); Zhong et al. (2022); Ryu et al. (2024a), and given that our goal is to control the summary rather than make it similar to the reference, we still include ROUGE and BERTScore Zhang et al. (2020) in our evaluation to provide clearer assessment. Our framework exhibits higher scores than other approaches, demonstrating through various evaluation metrics that GTE not only enhances controllability, but also improves overall summary quality.
@@ 8 Mixed attributes
We extend our evaluation to assess whether LLMs could precisely handle mixed-attribute control. While models manage to control linguistic attributes but struggle with numerical attributes. Simultaneous control over all attributes remains challenging for all iterative methods, including GTE. Our GTE framework guides LLMs to identify the causes of their errors and regenerate summaries by incorporating this feedback. However, in a mixed attribute setting, the model needs to handle multiple SAI and GS for each attribute simultaneously, increasing the cognitive load and making precise control of all attributes more challenging. Therefore, unlike single-attribute evaluation, which assesses whether attributes are precisely controlled, we evaluate mixed-attribute control by measuring errors using root mean squared error (rMSE). We compare the error between the attributes of the generated summary and the requested values, providing a more flexible evaluation of attribute control.
@@
Discovering the challenges in precisely controlling all attributes simultaneously, we introduce a sequential planning strategy, min-planning, which gradually adjusts attributes from the ill-controlled with the initial draft using GTE. Figure 3 shows the results comparing single attribute control with iterations to mixed-attribute control using min-planning on the MACSumDoc dataset. We refer to the initial summary in the mixed-attribute control as the mixed-draft. The min-planning method shows a modest improvement in controlling both attributes compared to the mixed-draft. Attributes are still not fully controlled compared to single-attribute models, highlighting the challenge of balancing multiple attributes. We anticipate that modifying one attribute often disrupts previously adjusted attributes due to correlations. For example, even if length is adjusted first, it may still change when extractiveness is controlled. Additionally, min-planning only adjusts each attribute once without iterations, which may account for its failure to fully control the attributes. A single refinement is often insufficient to control all attributes, whereas GTE iteratively regenerates the summary until the attribute is successfully adjusted in single-attribute control. Exploring ways for LLMs to control multiple attributes simultaneously would be promising future work.
Figure 3: Performance in mixed-attribute.
@@ 9 Conclusion
In this work, we systematically explore the controllability of LLMs. To this end, we revisit the measurement of various attributes. We evaluate the controllability of LLMs via iterative assessment, finding that LLMs struggle more with numeric attribute control than linguistic attributes. To address this, we propose a guide-to-explain (GTE) approach, where the model is guided to explain its misalignments through attribute identification and then grounds this explanation to generate better-controlled summaries in subsequent iterations. GTE enables LLMs to control challenging numerical attributes with lower failure and fewer iterations. Furthermore, we demonstrate the high quality of controlled summaries using various generic summarization evaluation metrics.
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00217286) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH)).

### v3

@@ 5 Guide-to-Explain (GTE)
Llama3-8B-GTE 12.63% / 3.52 11.63% / 2.53 0.00% / 1.46 11.70% / 3.26 26.40% / 3.08 10.92% / 2.26 13.18% / 3.85 14.99% / 2.80 25.56% / 0.91
Llama3-70B-Iter 54.82% / 8.44 37.21% / 7.47 2.70% / 3.78 49.91% / 8.05 18.40% / 6.58 54.61% / 10.42 67.44% / 12.00 49.36% / 8.24 0.00% / 0.24
Llama3-70B-SAI 26.55% / 6.57 18.60% / 7.81 0.00% / 1.86 24.14% / 6.52 4.80% / 5.42 2.73% / 3.81 10.85% / 4.84 5.12% / 4.39 0.00% / 0.10
Llama3-70B-GTE 0.21% / 3.28 0.00% / 2.83 0.00% / 1.50 0.18% / 3.22 0.00% / 1.10 0.00% / 1.61 2.32% / 3.14 0.55% / 1.90 0.00% / 0.01
GPT-3.5-Iter 45.18% / 9.80 60.47% / 0.00 94.59% / 0.00 49.73% / 9.80 53.60% / 0.00 80.89% / 0.00 88.37% / 0.00 76.42% / 0.00 0.00% / 0.00
GPT-3.5-GTE 17.56% / 3.86 51.16% / 5.00 67.57% / 4.00 23.58% / 3.90 5.60% / 4.63 44.03% / 6.62 78.29% / 7.00 43.33% / 5.95 0.00% / 0.00
GPT-4o-Iter 34.69% / 6.77 55.81% / 0.00 78.38% / 3.00 39.31% / 6.63 72.00% / 0.00 64.85% / 0.00 79.07% / 0.00 69.84% / 0.00 0.38% / 0.02
GPT-4o-SAI 35.12% / 5.50 48.84% / 15.50 62.16% / 6.00 38.03% / 6.13 60.00% / 8.79 61.09% / 9.40 78.29% / 2.00 64.90% / 8.60 0.00% / 0.04
GPT-4o-GTE 0.00% / 2.76 0.00% / 4.70 0.00% / 2.03 0.00% / 2.87 0.00% / 1.20 0.00% / 1.21 0.00% / 1.96 0.00% / 1.42 0.00% / 0.02
[…]
Model Extractiveness ($\downarrow$ / $\downarrow$) Length ($\downarrow$ / $\downarrow$) Topic ($\downarrow$ / $\downarrow$) Speaker ($\downarrow$ / $\downarrow$)
@@ 5.1 Step-by-step attribute identification
We first instruct the LLM to generate an initial summary $s^{\prime}$ that reflects the specified attribute. If the LLM fails to control the attributes accurately, we provide step-by-step attribute identification (SAI) to guide the model on how to adjust them with tools. Since LLMs struggle with measuring numerical attributes such as extractiveness or the length ratio, we explicitly instruct the model on how to measure each attribute step by step, enabling it to revise its previously generated summary more precisely.
@@ 5.2 Self-explanation guidance
After the identification phase, we provide self-explanation guidance (SEG) to the model, guiding the model to explain why it initially failed to control the attributes. This mirrors how humans solve complex problems by reviewing their mistakes to improve future responses. Building on this, in the next iteration, the document ($d$), initial instruction ($i$), and previously generated summary ($s^{\prime}$) are provided as inputs, along with SAI and SEG. Although LLMs are known to struggle with number-related tasks Akhtar et al. (2023); Imani et al. (2023), our guidance helps the model effectively control numerical attributes by self-explaining its miscalculations before generating summaries, especially when combined with the step-by-step attribute identification phase. We introduce GTE as a framework that integrates step-by-step attribute identification and self-explanation guidance.
@@ 5.3 Overall process
Figure 3 illustrates in detail how the GTE framework operates. By receiving [$d$; $i$, $s^{\prime}$; SAI; SEG] as input, the model first reflects on the reasons for its initial error before generating a revised summary. If the revised summary still fails to satisfy the attributes, GTE repeats the process until the model generates an attribute-aligned summary. See Appendix B for the detailed prompts.
@@ 6 Experimental setup
We evaluate the controllability of various LLMs, including Phi-3-medium Abdin et al. (2024), the Llama3 series Dubey et al. (2024), and the GPT series Brown (2020); Achiam et al. (2023). To analyze model performance by size, we utilize both the 8B 111meta-llama/Meta-Llama-3-8B-Instruct and quantized 70B versions222casperhansen/llama-3-70b-instruct-awq of Llama3, as well as GPT-3.5 and GPT-4o. We use BERTScore Zhang et al. (2020) to measure embedding similarity. We used two datasets for our experiments: MACSumDoc and MACSumDial Zhang et al. (2023b), which comprise committee meeting transcripts and news content, respectively. Both datasets are designed for mixed-attribute summarization that controls multiple attributes simultaneously. Notably, only MACSumDial include speaker attribute. Since we evaluate LLM performance on individual attributes, we use attributes separately.
@@ 7 Results and Discussions
Figure 4: The graphs show how the length ratio changes for each iteration. The intensity of the distribution color is proportional to the number of data points, and the markers represent the average values. The red line indicates the target length, with values of 7.5%, 20%, and 32.5% from left to right.
We denote the naive iteration approach, which repeatedly adjusts attributes, as Iter. The strategy that provides step-by-step attribute identification is denoted as SAI—a tool-augmented, stronger version of self-correction that provides the correct attribute values, since LLMs struggle to measure summary attributes on their own. As shown in Table 2, our GTE demonstrates remarkably lower failure rates and requires fewer iterations when adjusting summaries across all attributes, including challenging numerical attributes in MACSumDoc. Surprisingly, GTE reduced the failure rate to nearly 0% when applied to Llama3-70B and GPT-4o, successfully controlling both extractiveness and length within just 1–3 iterations. For smaller models such as Phi-3-medium and Llama3-8B, which initially exhibited high failure rates, our approach significantly reduced those rates, demonstrating its effectiveness across different model scales. In particular, for long length—the most challenging attribute—GTE achieved a remarkably low failure rate of just 2.32% within an average of 3.14 iterations.
LLMs encounter greater difficulty with the MACSumDial dataset (Table 3). The dataset, derived from QMSum Zhong et al. (2021), consists of lengthy and diverse content from parliamentary and committee meetings, making it more challenging than the CNN-news-based MACSumDoc. Notably, length control proved to be the most challenging attribute in MACSumDial. This challenge is likely due to the dataset’s origin in long parliamentary transcripts, which makes it inherently difficult to generate summaries of a specific target length. While the model handled short-length summaries relatively well, difficulty increased significantly as the requested summary length grew. In fact, for long-length summaries, both GPT-4o-Iter and GPT-4o-SAI showed a 100% failure rate. However, GTE showed meaningfully improved length controllability. With GPT-4o, the average failure rate dropped below 50%. Notably, for normal-length summaries, the failure rate further reduced from over 90% to 9.82%. Regarding extractiveness, the Iter and SAI of GPT-4o exhibit relatively low iteration counts, as the models often exceed the maximum iteration. While their failure rates were close to 80%, GTE achieved a markedly lower failure rate at 17.28% with low iterations, demonstrating the effectiveness of our framework.
@@
To analyze how the attribute changes at each step, we track length adjustments per iteration (Figure 4). While all methods start with a similar distribution at the initial point, GTE consistently converges within approximately three iterations, maintaining a stable length adjustment pattern across samples. In contrast, Iter and SAI show inconsistent changes across samples, resulting in higher variance in length adjustments. This demonstrates that our method enables robust attribute control with fewer iterations, regardless of the data sample. For this experiment, we use Llama3-70B and randomly select 110 samples from the MACSumDoc test set.
@@
We observe that LLMs control linguistic attributes (topic and speaker) better than numerical attributes (extractiveness and length). This aligns with previous research in mathematical reasoning, where LLMs struggle with numerical features Akhtar et al. (2023), highlighting a broader challenge in precisely handling numerical constraints. From the perspective of the summarization task, extractiveness and length control the structure of the summary, whereas topic and speaker influence its content. Our findings suggest that LLMs are proficient at adjusting content to align with user preferences but struggle to generate summaries with specific structural constraints.
@@
Table 4: Among the iterative methods, GTE demonstrates both effective attribute control and noticeable improvements in summary quality.
@@
We evaluate the quality of generated summaries including UniEval Zhong et al. (2022) and QuestEval Scialom et al. (2021), as they correlate highly with human judgments and assess the overall quality of the summary itself. UniEval is a multi-dimensional evaluator that assesses coherence, consistency, fluency, and relevance of summaries. QuestEval measures precision and recall by leveraging a question-answering framework to compare the content between the source document and the generated summary without relying on the reference summary. Table 4 shows that our method’s summaries outperform across all UniEval dimensions and QuestEval, demonstrating effective attribute control while maintaining overall summary quality. While Iter and SAI generate misaligned summaries with lower relevance (i.e., how well a summary retains key information compared to the reference), GTE effectively aligns them, resulting in a substantial gain.
Although our primary objective is to control the summary rather than match the reference, we report ROUGE Lin (2004) and BERTScore Zhang et al. (2020) to ensure comprehensive evaluation, as they are representative metrics in text summarization. GTE achieves higher scores than other iterative approaches, demonstrating across various evaluation metrics that GTE not only enhances controllability but also improves overall summary quality.
@@ 8 Mixed attributes
We extend our evaluation to assess whether LLMs can precisely handle mixed-attribute control. While models manage to control linguistic attributes, they struggle with numerical attributes. Simultaneous control over all attributes remains challenging for all iterative methods, including GTE. In GTE, the model is guided to identify the causes of its errors and regenerate summaries by incorporating this feedback. However, in a mixed attribute setting, the model must process multiple instances of SAI and SEG for each attribute simultaneously, increasing the cognitive load and making precise control of all attributes more difficult. Therefore, unlike single-attribute evaluation—which assesses whether individual attributes are accurately controlled—we evaluate mixed-attribute control by measuring errors using mean absolute deviation (MAD). This approach compares the differences between the attributes of the generated summary and the requested values, providing a more flexible evaluation of attribute control.
@@
Recognizing the challenges in precisely controlling all attributes simultaneously, we introduce a sequential planning strategy, min-planning, which gradually adjusts attributes—starting with those that are most poorly controlled in the initial draft—using GTE. Figure 5 shows the results comparing single-attribute control with iterations to mixed-attribute control using min-planning on the MACSumDoc dataset. We refer to the initial summary in the mixed-attribute control setting as the mixed-draft. The min-planning method shows a modest improvement in controlling both attributes compared to the mixed-draft. However, attributes are still not fully controlled as in single-attribute models, highlighting the difficulty of balancing multiple attributes. We anticipate that modifying one attribute often disrupts previously adjusted attributes due to underlying correlations. For example, even if length is adjusted first, it may still change when extractiveness is subsequently controlled. Additionally, min-planning adjusts each attribute only once without iteration, which may explain its inability to fully control the attributes. A single refinement is often insufficient, whereas GTE iteratively regenerates the summary until the target attribute is successfully adjusted in single-attribute control. Exploring ways for LLMs to control multiple attributes simultaneously would be promising future work.
Figure 5: Correlations among attributes hinder LLMs’ ability to control them jointly in mixed-attribute setting.
@@ 9 Conclusion
In this work, we systematically explore the controllability of LLMs. To this end, we revisit the measurement of summary attributes. We evaluate the controllability of LLMs via iterative assessment and find that they struggle more with numerical attributes than linguistic ones. To address this, we propose GTE, a self-correction-based approach that first instructs the model to identify misaligned attributes and then guides it to explain its own error to generate better-controlled summaries in subsequent iterations. GTE enables LLMs to control challenging numerical attributes with lower failure rates and fewer iterations. Furthermore, we demonstrate the high quality of controlled summaries using various evaluation metrics.
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00217286) (45%); by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2025 (No: RS-2025-02413038, Development of an AI-Based Korean Diagnostic System for Efficient Korean Speaking Learning by Foreigners) (45%); and by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH)) (10%)